OpenStudy (anonymous):

Find the vertical asymptotes, if any, for the following function. f(x)=-3x^2/x^2+9x-22

OpenStudy (anonymous):

To find a vertical asymptote set the denominator to zero x^2+9x-22=0 (x-2)(x+11)=0 x=2 and x=-11 So a vertical asymptote occurs at x=2 and x=-11 :)